2. A solution containing 0.402 49 g of CoCl2 ∙ xH2O (a solid with an unknown number of waters of hydration) was exhaustively electrolyzed to deposit 0.09937 g of metallic cobalt on a platinum cathode by the reaction Co2+ + 2e-  Co(s). Calculate the number of moles of water per mole of cobalt in the reagent.

1
Expert's answer
2022-05-10T05:27:03-0400

Solution:

%Сo in CoCl2 is:

%Сo in CoCl2 = (Molar mass of Co / Molar mass of CoCl2) = (58.933 / 129.839) = 0.4539 or 45.39%

Therefore, the mass of CoCl2 in CoCl2∙xH2O is:

Mass of CoCl2 = (Mass of Co / %Сo in CoCl2) = (0.09937 g / 0.4539) = 0.21893 g


Mass of CoCl2∙xH2O = 0.40249 g

Therefore,

Mass of H2O = Mass of CoCl2∙xH2O − Mass of CoCl2 = 0.40249 g − 0.21893 g = 0.18356 g


Moles of CoCl2 = (0.21893 g CoCl2) × (1 mol CoCl2 / 129.839 g CoCl2) = 0.001686 mol CoCl2

Moles of H2O = (0.18356 g H2O) × (1 mol H2O / 18.0153 H2O) = 0.010189 mol H2O


Divide the moles by the smallest of the results:

CoCl2 : 0.001686 / 0.001686 = 1.00

H2O : 0.010189 / 0.001686 = 6.04

Thus, the formula of CoCl2∙xH2O is CoCl2∙6H2O (6 moles of water per mole of cobalt)


Answer: 6 moles of water per mole of cobalt in the reagent